OpenStudy (anonymous):

find the limx→1(sinΠx)/(x−1) t=x-1 *without using l'hopital rule answer should be -Π but I don't know how to do without l'hopital rule

myininaya (myininaya):

is that a pi symbol?

OpenStudy (anonymous):

yup

myininaya (myininaya):

i will post something in just a sec k?

OpenStudy (anonymous):

okay, no biggie, and dont use l'hopital rule, make it as clear as you think I can catch up

myininaya (myininaya):

check that attachment out

myininaya (myininaya):

let me know if you have any questions

myininaya (myininaya):

remember as g->0, sing/g->1

myininaya (myininaya):

hey polpak can you prove the sin(x+y)=sinxcosy+sinycosx off the top of your head

OpenStudy (anonymous):

I can, but I usually do it with euler's stuff..

myininaya (myininaya):

hey heart, so you like totally got it?

myininaya (myininaya):

oh don't worry about showing me i will look up later

myininaya (myininaya):

i just forgot how to prove it

OpenStudy (anonymous):

\[sin(x+y) = Im( e^{(x+y)i})\]\[= Im (e^{xi} \cdot e^{yi})\]\[e^{xi} \cdot e^{yi} = (cos(x) + isin(x))(cos(y) + isin(y))\]\[= cos(x)cos(y) - sin(x)sin(y) \]\[+ i(cos(x)sin(y) + sin(x)cos(y)) \] \[\implies Im (e^{xi} \cdot e^{yi}) =cos(x)sin(y) + sin(x)cos(y) = sin(x+y)\]

OpenStudy (anonymous):

Where Im(a) is the imaginary part of a.

myininaya (myininaya):

thats real pretty

OpenStudy (anonymous):

A lot of the trig identities become really pretty when you use eulier's formula: \[e^{\theta i} = cos(\theta) + isin(\theta)\]

OpenStudy (anonymous):

And for free you get the identity for cos(x+y) too.. It's just the real part of the same thing so you get \[cos(x+y) = cos(x)cos(y) - sin(x)sin(y)\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

What's funny?

myininaya (myininaya):

for free

OpenStudy (anonymous):

Well sorta free since you had to find it to get the imaginary part ;p

myininaya (myininaya):

i get a discount

myininaya (myininaya):

thanks polpak

OpenStudy (anonymous):

Of course! /tips lego hat